RC Meeting, Monday, October 04 , 2021

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 10 am, remote

  • Bluejeans remote connection information:
 To join the Meeting:
 https://bluejeans.com/861439920
 To join via Room System:
 Video Conferencing System: bjn.vc -or-199.48.152.152
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Accelerator Status

  1. Yesterday, right after SHMS was rotated to 5.505 deg, Ops was in the process of doing ion chamber calibrations at the new angle, as it was found that the position at the dump shifted after spectrometer rotation which caused the ion chambers at the beam dump to trip.
  2. During the ion chamber calibrations, Accelerator South Linac "2K Cold Box Trip" which caused loss of beam for a few hours (~11:00 - 05:00) , but beam quickly recovered and has been very stable.

Hall C Status

  1. PION-LT production concluded with ~ 95% of PAC-approved pions collected, which exceeded the expectations we had a few days earlier.
  2. Yesterday early in the morning (~9:00-10:00), SHMS was rotated to 5.505 deg (a milestone for this experiment and for the Hall) --> SHMS Smallest Angle
  3. During the "Controlled Access" to rotate spectrometer, Brad installed a temporary replacement camera for HCCAM03 which had failed a few days ago --> HCCAM03 Broken, HCCAM03 replaced
  4. Today at 13:24, there was an HMS Dipole Interlock alarm. Contacted Amy Comer, but nothing could be done remotely. She contacted Joe Beaufait and he went to the Hall with RADCON to check the problem --> https://logbooks.jlab.org/entry/3917638

Current Run Plan

  1. Continue taking production data on LH2 target with 40 uA beam at new kinematics (SHMS: 8.035 GeV/c, 5.505 deg | HMS: -1.026 GeV/c, 47.07 deg)
    • As of run 12463 (ended @ 12:15), we have accumulated: 2,286.453 mC and 1,190 pions (11.9 % of PAC-approved statistics)
  2. NOTE: At the smallest SHMS angle (5.505 deg), the typical S1X Hodoscope Rates are ~2.4 MHz @ 40 uA beam --> SHMS S1X Scaler Rates, whereas at the previous SHMS angle (7.05 deg), the typical S1X Hodoscope rates were ~ 2.1 MHz @ 70 uA. I was informed by Owl shift crew that accelerator person had stopped by the CH to find out if we were planning to increase the beam current in the near future, since they expressed concerns when Hall A would start running. I informed Jay Benesch and he mentioned that this should not be an issue as the accelerator had the capacity to hold up both Halls.

Current Issues

  1. Target ladder camera was identified to be bad: Joe Beaufait will purchase a replacement in the next funding cycle ( PENDING )
    • Still need to follow-up with Joe about when the purchase can be done.
  2. HCCAM03 Hall Camera has failed and needs replacement --> HCCAM03 Broken ( CLEARED )
    • Brad accessed the Hall during "Controlled Access" and replaced the broken HCCAM03 with the SHMS Hut camera for the bubbler,
      as the spare was unfortunately broken as well. See log-entry --> https://logbooks.jlab.org/entry/3916985

Special Requests

  1. Contact Ellen Becker if there is an opportunity to access the hall (longer than 4 hours). This is for diagnosing the SHMS power supply. ( PENDING )